THERE IS 3 BLACK BALL 2 YELLOW AND 4 RED BALLS WHATS THE PROBABILITY THAT I'M GETTING A BLACK BALL
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)
Answer by checkley77(12844)

3/9=1/3=.3333=33.33% probability of getting a black ball.